# A Bias–Variance Decomposition for Ensembles over Multiple Synthetic Datasets

### 06 Feb 2024 v1

@@ Appendix D Extra Results
D.1 Estimating Model and Synthetic Data Variances
@@
The result is an estimate of MV and SDV for each test point. We plot the mean over the test points. The whole experiment is repeated 3 times, with different train-test splits. The datasets, train-test splits, and downstream predictors are the same as in the other experiments, described in Appendix C.
The results are in Figure 13. MV depends mostly on the downstream predictor, while SDV also depends on the synthetic data generator. We also confirm that decision trees and 1-NN have much higher variance than the other models, and linear, ridge and logistic regression have a very low variance.
@@
Figure 13: Estimating the MV and SDV terms from the decomposition. Decision trees have high variances on all datasets, while linear, ridge and logistic regression have low variances. MV depends mostly on the predictor, while SDV depends on both the predictor and synthetic data generation algorithm. The points are the averages of estimated MV and SDV, averaged over the test data, from 3 repeats with different train-test splits. We excluded some test data points that had extremely large variance estimates ($\geq 10^{6}$) for linear regression on the ACS 2018 dataset. The predictors are decision tree (DT), nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), linear regression (LR), ridge regression (RR) and logistic regression (LogR). The synthetic data generators are DDPM and synthpop (SP-P).

### 23 May 2024 v2

@@ E.3 Result Tables
Table S12: Table of results from predicting Brier score on the German Credit dataset. The numbers are the mean measured and predicted Brier scores $\pm$ standard deviations from 3 repeats.
m 1 2 4 8 16 32
Downstream Generator Predicted / Measured
1-NN DDPM Measured 0.41 $\pm$ 0.027 0.33 $\pm$ 0.024 0.32 $\pm$ 0.032 0.30 $\pm$ 0.044 0.29 $\pm$ 0.036 0.28 $\pm$ 0.038
Predicted 0.41 $\pm$ 0.027 0.33 $\pm$ 0.024 0.30 $\pm$ 0.022 0.28 $\pm$ 0.021 0.27 $\pm$ 0.021 0.27 $\pm$ 0.021
SP-P Measured 0.37 $\pm$ 0.033 0.28 $\pm$ 0.029 0.25 $\pm$ 0.030 0.23 $\pm$ 0.035 0.21 $\pm$ 0.035 0.20 $\pm$ 0.033
Predicted 0.37 $\pm$ 0.033 0.28 $\pm$ 0.029 0.24 $\pm$ 0.037 0.22 $\pm$ 0.043 0.21 $\pm$ 0.046 0.20 $\pm$ 0.047
5-NN DDPM Measured 0.24 $\pm$ 0.020 0.24 $\pm$ 0.015 0.23 $\pm$ 0.019 0.23 $\pm$ 0.024 0.23 $\pm$ 0.025 0.23 $\pm$ 0.025
Predicted 0.24 $\pm$ 0.020 0.24 $\pm$ 0.015 0.23 $\pm$ 0.017 0.23 $\pm$ 0.020 0.23 $\pm$ 0.021 0.23 $\pm$ 0.022
SP-P Measured 0.23 $\pm$ 0.009 0.21 $\pm$ 0.018 0.20 $\pm$ 0.017 0.19 $\pm$ 0.018 0.19 $\pm$ 0.015 0.19 $\pm$ 0.018
Predicted 0.23 $\pm$ 0.009 0.21 $\pm$ 0.018 0.20 $\pm$ 0.024 0.20 $\pm$ 0.026 0.19 $\pm$ 0.028 0.19 $\pm$ 0.029
Decision Tree DDPM Measured 0.44 $\pm$ 0.043 0.32 $\pm$ 0.040 0.28 $\pm$ 0.038 0.27 $\pm$ 0.040 0.26 $\pm$ 0.037 0.25 $\pm$ 0.040
Predicted 0.44 $\pm$ 0.043 0.32 $\pm$ 0.040 0.26 $\pm$ 0.039 0.23 $\pm$ 0.039 0.21 $\pm$ 0.038 0.20 $\pm$ 0.038
SP-P Measured 0.33 $\pm$ 0.033 0.25 $\pm$ 0.016 0.23 $\pm$ 0.027 0.21 $\pm$ 0.031 0.19 $\pm$ 0.026 0.18 $\pm$ 0.022
Predicted 0.33 $\pm$ 0.033 0.25 $\pm$ 0.016 0.22 $\pm$ 0.007 0.20 $\pm$ 0.003 0.19 $\pm$ 0.001 0.18 $\pm$ 0.001
Random Forest DDPM Measured 0.22 $\pm$ 0.013 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009
Predicted 0.22 $\pm$ 0.013 0.21 $\pm$ 0.008 0.21 $\pm$ 0.006 0.21 $\pm$ 0.005 0.21 $\pm$ 0.004 0.21 $\pm$ 0.004
SP-P Measured 0.19 $\pm$ 0.005 0.18 $\pm$ 0.010 0.17 $\pm$ 0.008 0.17 $\pm$ 0.007 0.17 $\pm$ 0.006 0.17 $\pm$ 0.007
Predicted 0.19 $\pm$ 0.005 0.18 $\pm$ 0.010 0.17 $\pm$ 0.012 0.17 $\pm$ 0.014 0.17 $\pm$ 0.014 0.17 $\pm$ 0.015
MLP DDPM Measured 0.33 $\pm$ 0.022 0.29 $\pm$ 0.019 0.28 $\pm$ 0.025 0.27 $\pm$ 0.032 0.26 $\pm$ 0.030 0.26 $\pm$ 0.026
Predicted 0.33 $\pm$ 0.022 0.29 $\pm$ 0.019 0.28 $\pm$ 0.023 0.27 $\pm$ 0.026 0.26 $\pm$ 0.027 0.26 $\pm$ 0.028
SP-P Measured 0.27 $\pm$ 0.033 0.23 $\pm$ 0.027 0.19 $\pm$ 0.022 0.18 $\pm$ 0.019 0.18 $\pm$ 0.020 0.17 $\pm$ 0.018
Predicted 0.27 $\pm$ 0.033 0.23 $\pm$ 0.027 0.21 $\pm$ 0.044 0.20 $\pm$ 0.054 0.19 $\pm$ 0.059 0.19 $\pm$ 0.061
Gradient Boosting DDPM Measured 0.23 $\pm$ 0.019 0.22 $\pm$ 0.011 0.22 $\pm$ 0.008 0.22 $\pm$ 0.008 0.22 $\pm$ 0.007 0.21 $\pm$ 0.010
Predicted 0.23 $\pm$ 0.019 0.22 $\pm$ 0.011 0.22 $\pm$ 0.008 0.22 $\pm$ 0.007 0.22 $\pm$ 0.007 0.22 $\pm$ 0.007
SP-P Measured 0.20 $\pm$ 0.016 0.18 $\pm$ 0.016 0.17 $\pm$ 0.011 0.17 $\pm$ 0.012 0.17 $\pm$ 0.009 0.16 $\pm$ 0.010
Predicted 0.20 $\pm$ 0.016 0.18 $\pm$ 0.016 0.17 $\pm$ 0.018 0.17 $\pm$ 0.020 0.16 $\pm$ 0.021 0.16 $\pm$ 0.022
SVM DDPM Measured 0.21 $\pm$ 0.009 0.21 $\pm$ 0.008 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009 0.21 $\pm$ 0.009
Predicted 0.21 $\pm$ 0.009 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008 0.21 $\pm$ 0.008
SP-P Measured 0.20 $\pm$ 0.007 0.18 $\pm$ 0.012 0.17 $\pm$ 0.011 0.17 $\pm$ 0.010 0.17 $\pm$ 0.007 0.17 $\pm$ 0.008
Predicted 0.20 $\pm$ 0.007 0.18 $\pm$ 0.012 0.17 $\pm$ 0.016 0.17 $\pm$ 0.018 0.17 $\pm$ 0.019 0.17 $\pm$ 0.020
Logistic Regression DDPM Measured 0.23 $\pm$ 0.015 0.22 $\pm$ 0.006 0.21 $\pm$ 0.007 0.21 $\pm$ 0.011 0.21 $\pm$ 0.011 0.21 $\pm$ 0.011
Predicted 0.23 $\pm$ 0.015 0.22 $\pm$ 0.006 0.21 $\pm$ 0.003 0.21 $\pm$ 0.003 0.21 $\pm$ 0.004 0.21 $\pm$ 0.004
SP-P Measured 0.20 $\pm$ 0.003 0.18 $\pm$ 0.010 0.17 $\pm$ 0.012 0.17 $\pm$ 0.009 0.17 $\pm$ 0.008 0.16 $\pm$ 0.008
Predicted 0.20 $\pm$ 0.003 0.18 $\pm$ 0.010 0.17 $\pm$ 0.017 0.17 $\pm$ 0.020 0.16 $\pm$ 0.021 0.16 $\pm$ 0.022
Table S13: Table of results from the DP experiment. The numbers are the mean Brier score $\pm$ the standard deviation from 3 repeats. The privacy parameters are $\epsilon=1.5$, $\delta=n^{-2}\approx 4.7\cdot 10^{-7}$.
m 1 2 4 8 16 32
Downstream Generator
1-NN AIM - Log Prob. Avg. 0.23 $\pm$ 0.012 0.17 $\pm$ 0.003 0.16 $\pm$ 0.002 0.16 $\pm$ 0.011 0.15 $\pm$ 0.006 0.15 $\pm$ 0.001
AIM - Prob. Avg. 0.23 $\pm$ 0.012 0.17 $\pm$ 0.003 0.14 $\pm$ 0.003 0.13 $\pm$ 0.004 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.25 $\pm$ 0.008 0.20 $\pm$ 0.020 0.19 $\pm$ 0.005 0.17 $\pm$ 0.006 0.16 $\pm$ 0.005 0.16 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.25 $\pm$ 0.008 0.20 $\pm$ 0.020 0.16 $\pm$ 0.003 0.14 $\pm$ 0.005 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003
5-NN AIM - Log Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.002 0.14 $\pm$ 0.002 0.14 $\pm$ 0.002 0.13 $\pm$ 0.004 0.13 $\pm$ 0.001
AIM - Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.16 $\pm$ 0.007 0.15 $\pm$ 0.004 0.15 $\pm$ 0.008 0.15 $\pm$ 0.008 0.14 $\pm$ 0.003 0.14 $\pm$ 0.003
NAPSU-MQ - Prob. Avg. 0.16 $\pm$ 0.007 0.14 $\pm$ 0.005 0.13 $\pm$ 0.005 0.13 $\pm$ 0.005 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
Decision Tree AIM - Log Prob. Avg. 0.13 $\pm$ 0.003 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.13 $\pm$ 0.003 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.16 $\pm$ 0.005 0.15 $\pm$ 0.004 0.15 $\pm$ 0.005 0.14 $\pm$ 0.004 0.14 $\pm$ 0.004 0.14 $\pm$ 0.003
NAPSU-MQ - Prob. Avg. 0.16 $\pm$ 0.005 0.14 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
Random Forest AIM - Log Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.14 $\pm$ 0.007 0.14 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.14 $\pm$ 0.007 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003
MLP AIM - Log Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.14 $\pm$ 0.006 0.13 $\pm$ 0.004 0.13 $\pm$ 0.005 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.14 $\pm$ 0.006 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.12 $\pm$ 0.004
Gradient Boosting AIM - Log Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.11 $\pm$ 0.002 0.11 $\pm$ 0.001 0.11 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.003 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004
SVM AIM - Log Prob. Avg. 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.001 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002
AIM - Prob. Avg. 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002 0.13 $\pm$ 0.001 0.13 $\pm$ 0.002 0.13 $\pm$ 0.002
NAPSU-MQ - Log Prob. Avg. 0.14 $\pm$ 0.005 0.14 $\pm$ 0.006 0.14 $\pm$ 0.005 0.14 $\pm$ 0.004 0.13 $\pm$ 0.005 0.13 $\pm$ 0.005
NAPSU-MQ - Prob. Avg. 0.14 $\pm$ 0.005 0.14 $\pm$ 0.006 0.14 $\pm$ 0.005 0.14 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.005
Logistic Regression AIM - Log Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001
AIM - Prob. Avg. 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.002 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001 0.12 $\pm$ 0.001
NAPSU-MQ - Log Prob. Avg. 0.13 $\pm$ 0.005 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
NAPSU-MQ - Prob. Avg. 0.13 $\pm$ 0.005 0.13 $\pm$ 0.003 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004 0.13 $\pm$ 0.004
E.4 Estimating Model and Synthetic Data Variances
@@
The result is an estimate of MV and SDV for each test point. We plot the mean over the test points. The whole experiment is repeated 3 times, with different train-test splits. The datasets, train-test splits, and downstream predictors are the same as in the other experiments, described in Appendix D.
The results are in Figure S11. MV depends mostly on the downstream predictor, while SDV also depends on the synthetic data generator. We also confirm that decision trees and 1-NN have much higher variance than the other models, and linear, ridge and logistic regression have a very low variance.
@@
Figure S11: Estimating the MV and SDV terms from the decomposition. Decision trees have high variances on all datasets, while linear, ridge and logistic regression have low variances. MV depends mostly on the predictor, while SDV depends on both the predictor and synthetic data generation algorithm. The points are the averages of estimated MV and SDV, averaged over the test data, from 3 repeats with different train-test splits. We excluded some test data points that had extremely large variance estimates ($\geq 10^{6}$) for linear regression on the ACS 2018 dataset. The predictors are decision tree (DT), nearest neighbours with 1 or 5 neighbours (1-NN and 5-NN), random forest (RF), a multilayer perceptron (MLP), gradient boosted trees (GB), a support vector machine (SVM), linear regression (LR), ridge regression (RR) and logistic regression (LogR). The synthetic data generators are DDPM and synthpop (SP-P).